OpenStudy (anonymous):

I need some help solving a system of equations...Can someone help me please?

OpenStudy (anonymous):

2x+2y+z=-5 3x+4y+2z=0 x+3y+2z=1

OpenStudy (jack1):

ok... so what have you done so far?

OpenStudy (anonymous):

I think I figured it out...is the answer x=-10, y=19, z=-23??

OpenStudy (jack1):

yep, perfect

OpenStudy (anonymous):

yay!! Can you check a few of my oher questions?

OpenStudy (jack1):

sure, just close this one down and open a new q, yeah?

OpenStudy (anonymous):

okay!